# Topic-FlipRAG: Topic-Orientated Adversarial Opinion Manipulation Attacks to Retrieval-Augmented Generation Models

### v2

@@ 4.3.2. Gradient-Based Optimization
Gradient Optimization: During each optimization step, we compute and combine gradients for all queries in $\mathbf{Q}$. Additionally, a word enhancement argument is introduced, which incorporates query-specific keywords and lexical distributions from the model’s vocabulary. This guided selection of candidate tokens captures semantic overlaps among multiple queries and enhances overall trigger relevance.
@@
Beam Search: In the beam search phase, we evaluate candidate triggers—along with doc-specific augmentation—to select those that maximize combined similarity scores across the entire query set. This ensures the chosen trigger reliably improves the target passage’s ranking impact when integrated with $\text{doc}_{\text{know}}$.
@@
Doc-Specific Augmentation: At each iteration of gradient calculation and beam search, the trigger’s intermediate soft representation is evaluated only when concatenated with $\text{doc}_{\text{know}}$ and considered jointly with all queries. Although this increases computational overhead, it ensures tight alignment with the target document’s context, yielding more pronounced improvements in ranking performance.
4.3.3. Final Output: $\text{doc}_{\text{adv}}$ and RAG Opinion Influence
@@
After convergence, the adversarial trigger $T$ is appended to $\text{doc}_{\text{know}}$ to form the final adversarial document:
$\text{doc}_{\text{adv}}=[\text{doc}_{\text{know}},T].$
@@
This $\text{doc}_{\text{adv}}$ is then injected into the RAG system’s corpus $D^{\prime}$, significantly increasing its likelihood of being retrieved among the top-$k$ documents for each query $q\in Q$. Consequently, $\text{doc}_{\text{adv}}$ frequently appears in retrieval results, leading the LLM to generate outputs that align with the desired stance $S_{t}$. This process effectively manipulates the overall stance of the RAG system by ensuring that the augmented document is both prominently retrieved and influences the generated opinions towards $S_{t}$.
5. Experiments
5.1. Datasets
MSMARCO. We utilized the MS MARCO Passage Ranking dataset as the data source to evaluate the ability of our method to improve document rankings under more challenging topic-query tasks. Specifically, we assessed whether our method could significantly enhance the ranking of documents by the retrieval model within a RAG system.
To construct topic-lists for evaluation, we applied a K-means clustering algorithm to group similar queries, forming topics that each contained a series of related queries. To further evaluate the performance of our method under extreme topic-query scenarios, we applied an intra-topic similarity filtering process. Only topics with queries exhibiting high semantic diversity and containing a sufficient number of queries were retained.
This process resulted in 29 topics, with each topic containing an average of 22.28 queries. The average similarity score within each topic was approximately 0.5, indicating sufficient diversity among queries to ensure a rigorous evaluation. This curated dataset enabled us to test the robustness of our method in handling highly diverse and challenging topic-query tasks within a RAG system.
PROCON. To conduct our experiments, we utilized controversial topic data scraped from the PROCON.ORG website. This dataset includes over 80 topics covering various domains, such as society, health, government, and education. Each topic is discussed from two stance labels {PRO (support), CON (oppose)}, with passages arguing from these perspectives.
To simulate real-world user interactions with a RAG system, we instructed a large language model (GPT-4o) to act as a user and generate 40 potential sub-queries for each topic. These sub-queries were designed to reflect the diverse questions and concerns users might raise when exploring a specific controversial topic.
After generating the sub-queries, we applied a similarity filtering process to ensure diversity by retaining only those with a similarity score below approximately 0.85. The filtering step effectively removed redundant queries while preserving a wide range of perspectives. As a result, the final set of topic-queries achieved an average similarity score of approximately 0.7, ensuring that the queries were sufficiently diverse yet semantically relevant. This process provided a robust and balanced sub-queries set for evaluation.
5.2. Experiment Details
(1) Black-box RAG. We represent the black-box RAG process as $\text{RAG}_{\text{black}}$. The RAG framework is Conversational RAG from LangChain. The LLMs adopted in RAG are the open-source models Meta-Llama-3.1-8B-Instruct (Llama3.1), Qwen-2.5-7B-Instruct(Qwen2.5). The system prompt and additional detailed descriptions are provided in Appendix A.1.
(2) Retrieval Model Specification. We benchmark three dominant dense retrievers—Contriever (Gao and Callan, 2021), DPR (Karpukhin et al., 2020), and ANCE (Xiong et al., 2020).By convention, we use dot product between the embedding vectors of questions(queries) and candidate documents as their similarity score $R$ in the ranking.
(3) Opinion classification. We use Qwen2.5-Instruct-72B as the opinion classifier. Qwen2.5-Instruct-72B, due to its large parameter size, is capable of accurately identifying and classifying opinions within text.
(4) Experimental parameters. In knowledge-guided attack process, we set the maximum editing distance $\epsilon$ to 0.2, the semantic similarity threshold $\lambda$ to 0.85, and the number of iterations $N$ to 5. For adversarial trigger generation, we used a beam size of 3, top-$k$ of 10, a batch size of 32, a temperature of 1.0, a learning rate of 0.005, and a sequence length of 10. In RAGblack, $k$ (the number of retrieved documents) is set to 3, with the LLM temperature also fixed at 1.0 to mirror real-world conditions.
(5) Poisoned Target. For the PROCON dataset, to investigate the manipulation performance under more challenging conditions, we performed relevance ranking for the documents with respect to each topic-query set $Q$ and the target stance $S_{t}$ From the ranked list, we selected the last five documents (i.e., those with the lowest relevance) as the target poisoned documents. For the MS MARCO dataset, we utilized the top-1000 relevance-ranked passage list for each topic-query set. From this list, we selected the passage with the lowest average rank as the target passage. This approach ensures that the evaluation focuses on passages that are least relevant to the target queries, thus providing a more rigorous benchmark for the proposed method.
(6) Experimental environment. All our experiments are conducted in Python 3.8 environment and run on a NVIDIA DGX H100 GPU.
5.3. Research Questions
@@
We propose four research questions to evaluate the effectiveness of our method in the topic-queries task, focusing on black-box NRM attacks and opinion manipulation to RAGs.
@@
RQ1: Can Topic-FlipRAG significantly enhance the rankings of target documents in the RAG retriever for topic-queries?
@@
RQ4: How robust does Topic-FlipRAG against exisiting mitigation mechanism?
5.4. Baseline Settings
@@
To assess the effectiveness of our proposed method, we compare it against adversarial attack baselines designed for black-box, topic-oriented RAG scenarios, ensuring minimal modifications to the original documents. We exclude BadRAG(Xue et al., 2024), a backdoor RAG attack limited to white-box scenarios, and topic-IR-attack(Liu et al., 2023c), as its incomplete implementation prevents reliable replication. For the selected baseline methods, we adapted them to meet the requirements of our task while preserving their core components. A brief overview of the baseline methods is provided below, with detailed descriptions available in Appendix A.2.
@@
PoisonedRAG. Zou et al.(Zou et al., 2024) propose an approach adaptable to both black-box and white-box settings. For our task, we employ its black-box strategy by inserting a randomly chosen query from the topic-queries set $Q$ at the beginning of each document.
@@
PAT. This gradient-based adversarial retrieval attack uses a pairwise loss function to generate triggers that meet fluency and coherence constraints. We adapt PAT to produce triggers $T_{\text{pat}}$ for target documents within the topic-queries set, evaluating their effectiveness under black-box conditions.
@@
Collision. This method generates adversarial paragraphs (collisions) via gradient-based optimization to produce content semantically aligned with the target query. In a topic-queries context, we create collisions for the entire topic-queries set and examine their transfer performance on black-box RAG retrievers.
@@
These baseline methods provide benchmarks for comparing the efficacy of our approach in a fully black-box, topic-oriented RAG attack scenario.
5.5. Evaluation Metrics
@@
Ranking Attack Success Rate (RASR). RASR captures how frequently target documents are successfully boosted in each query’s ranking. Higher values indicate greater attack effectiveness.
Boost Rank (BRank). BRank is the average rank improvement for all target documents under each query. A target document contributes negatively if its rank is unintentionally lowered.
For RQ2, we employ Average Stance Variation (ASV) to assess how significantly our opinion manipulation influences the LLM’s responses in a black-box RAG. To address the natural variability of controversial topics and the inherent instability of large language models, we also propose Real Adjusted ASV ($\Delta$-ASV).
Average Stance Variation (ASV). ASV is defined as the absolute difference between the manipulated opinion score and the original opinion score assigned to an LLM response (0 = opposing, 1 = neutral, 2 = supporting). A higher ASV signifies a more pronounced shift in polarity and hence greater manipulation effectiveness.
[…]
6. Experimental Results Analysis
6.1. RQ1: Can our method significantly enhance the rankings of target documents in the RAG retriever for topic-queries?
To explore RQ1, we first conducted white-box IR manipulation experiments on MS MARCO and PROCON datasets. The results of these experiments are presented in Table 1 and Table 2, respectively.
MS MARCO Dataset Validation. The results demonstrate that Topic-FlipRAG significantly outperforms all other baselines across all evaluation metrics, even under challenging conditions where the MS MARCO dataset was configured with reduced internal similarity among topic-queries to simulate more extreme scenarios. On this dataset (Table 1), Topic-FlipRAG achieves an almost perfect Retrieval Attack Success Rate (RASR) of 99.29%, substantially higher than Poisoned-RAG (64.39%), Collision (93.19%), and PAT (89.63%). Furthermore, it achieves the highest B-rank (521.30) and superior coverage of top-ranked documents, with 18.35% for top-50% and 79.49% for top-500% metrics. These results highlight Topic-FlipRAG’s ability to effectively manipulate rankings across the entire topic-queries set, achieving significant ranking enhancements for each query while simultaneously targeting multiple queries with high success rates.
Table 1. White-box IR attack performance (RASR, Top-50, and Top-500 are reported in %) on MS MARCO dataset. Bold indicates the best attack performance.
@@ 6. Experimental Results Analysis
Table 2. White-box and Black-box IR attack results (%) on PROCON dataset. Bold indicates the best attack performance.
@@
PROCON Dataset Validation. On the PROCON dataset (Table 2), the results strongly confirm the superiority of Topic-FlipRAG. In the white-box setting, for both target stances (CON and PRO), Topic-FlipRAG achieves the highest RASR (0.79 for CON, 0.75 for PRO) and the largest top3-v (0.45 for CON, 0.37 for PRO). In contrast, Poisoned-RAG, which uses a simple query+ strategy, shows considerably lower effectiveness (RASR: 0.40 for CON, 0.33 for PRO), illustrating the limitations of its approach for topic-oriented attacks. While the gradient-based methods Collision and PAT match or exceed Poisoned-RAG in opinion manipulation, they still trail Topic-FlipRAG. This outcome highlights Topic-FlipRAG’s ability to conduct effective manipulation across entire sets of topic-queries.
To further evaluate our method’s practical effectiveness, we conducted black-box IR manipulation experiments on the PROCON dataset using three widely adopted dense retrievers (Contriever, ANCE, DPR). Owing to the inherent robustness of these retrievers, performance declined for all attack methods compared to the white-box setting. Nevertheless, Topic-FlipRAG displays the most consistent performance, retaining relatively high manipulation effectiveness across all models and metrics. For instance, under Contriever, Topic-FlipRAG achieves the highest RASR (0.52 for CON, 0.49 for PRO) and top3-v (0.26 for CON, 0.22 for PRO), notably outperforming Poisoned-RAG, Collision, and PAT. A similar pattern emerges for ANCE and DPR, where Topic-FlipRAG continues to hold a clear advantage.
Topic-FlipRAG’s superior performance in transfer attack settings can be attributed to its integration of the LLM’s general semantic knowledge and NRM gradient guidance, along with a design specifically tailored for the topic-queries context. This combination enables the method to generate adversarial triggers that are both generalizable and contextually precise, allowing it to adapt effectively to varying retriever architectures without relying on the target IR model’s parameters. In contrast, gradient-based methods like Collision and PAT depend heavily on gradient information from the target model, leading to substantial performance degradation in transfer attack scenarios. For instance, Collision achieves a maximum RASR of only 0.21 on Contriever for both CON and PRO stances, while PAT performs even worse, with RASR values of 0.16 (CON) and 0.13 (PRO).
Table 3. Topic-Oriented RAG attack results on PROCON dataset, including ASV on different domains (Health & Environment, Government & Politics, Education, and Society & Culture), average ASV, and $\Delta$ASV. Bold shows the best performances.
Method Target Stance Contriever + qwen-2.5-8B-instruct Contriever + llama3.1-8B-instruct
H&E G&P EDU S&C avg. ASV $\Delta$ ASV H&E G&P EDU S&C avg. ASV $\Delta$ ASV
@@
6.2. RQ2: To what extent does Topic-FlipRAG affect the answers generated by the target RAG systems?
Figure 3. Case study of Topic-oriented Opinion Manipulation On Topic: ”Is Binge-Watching Good for You?”
Table 3 presents the results of different methods on RAG output opinion manipulation, comparing the performance of two retrieval-augmented generation (RAG) systems: contriever + qwen-2.5-8B-instruct and contriever + llama3.1-8B-instruct. Additionally, experiments were conducted using alternative retrievers such as ANCE and DPR, with results provided in Table8 and Table 9 (in Appendix).
To evaluate the manipulation effectiveness across different topic domains, the PROCON dataset was divided into four subsets: Health & Environment (H&E), Government & Politics (G&P), Education (EDU), and Society & Culture (S&C). Each reported avg-ASV reflects the average manipulated stance polarity (PRO or CON) of the RAG outputs across all domains, while avg-$\Delta$-ASV represents the manipulation effect, computed by subtracting the clean-state avg-ASV (clean) from the manipulated avg-ASV. Both ASV and $\Delta$-ASV range from 0 to 2, where higher values indicate stronger manipulation effects.
The results demonstrate that Topic-FlipRAG achieves significantly better manipulation across all domains compared to baseline methods, including PoisonedRAG. Specifically, Topic-FlipRAG achieves an avg-ASV of 0.49 for qwen-2.5 and 0.48 for llama3.1, outperforming all other methods. PoisonedRAG, the second most effective method, achieves avg-ASV values of 0.33 for qwen-2.5 and 0.35 for llama3.1, while other baselines such as Collision and PAT remain within the range of 0.28–0.31. In terms of manipulation effect, measured by avg-$\Delta$-ASV, Topic-FlipRAG also achieves the highest values: 0.26 for qwen-2.5 and 0.24 for llama3.1. These results highlight the robust manipulation capability of Topic-FlipRAG, surpassing both direct adversarial document injection methods like PoisonedRAG (0.09 for qwen-2.5 and 0.10 for llama3.1) and stance-optimization methods such as and PAT (0.06 for both LLMs).
Table 3 and the appendix results further reveal notable differences between LLMs and retrievers regarding susceptibility to topic-oriented opinion manipulation. Overall, llama3.1 exhibits higher avg-ASV and avg-$\Delta$-ASV across all baseline methods compared to qwen-2.5, indicating that llama3.1 is more vulnerable to adversarial documents promoted to the top-3 of the ranking. This suggests that llama3.1 tends to generate responses that are more aligned with adversarial stances in the context. Similarly, among retrievers, contriever-based RAG systems achieve higher avg-$\Delta$-ASV values compared to ANCE and DPR, as detailed in the appendix, aligning with RQ1 findings that contriever produces superior ranking distortion, particularly for top-3 results.
The manipulation effectiveness of Topic-FlipRAG also varies across different topic domains. For instance, for llama3.1, Society & Culture (S&C) questions are the most susceptible to opposing stance manipulation (avg-ASV = 0.55), while Education (EDU) questions are relatively resistant (avg-ASV = 0.41). Health & Environment (H&E) questions demonstrate greater difficulty in manipulating supporting stances (PRO) compared to opposing stances (CON). Both qwen-2.5 and llama3.1 display higher avg-ASV values for supporting stance manipulation in Government & Politics (G&P), while Education (EDU) emerges as the most challenging domain for stance manipulation across both LLMs.
In order to illustrate the manipulation results more vividly, we randomly selected the topic “Is Binge-Watching Good for You?” for a case study, with the findings presented in Figure 3. As demonstrated by the distribution variation of RAG response opinions in the figure,, Topic-FlipRAG can manipulate RAG-generated content across multiple queries within the specific topic, shifting the overall opinion distribution to match a specified stance. Initially, responses on this topic span a mix of supportive, opposing, and neutral perspectives; following manipulation, the distribution becomes predominantly supportive when the target stance is “Pro” (blue) or predominantly opposing when the target stance is “Con” (red).
Figure 3 further demonstrates Topic-FlipRAG’s impact on multiple sub-queries. For instance, in “Binge-watching: addictive or relaxing?”, the unmanipulated response presents a balanced view. However, once manipulated, it adopts a distinctly polarized stance—emphasizing the relaxing aspects under “Pro,” or highlighting detrimental, addictive effects under “Con.” Similarly, for “Opinions on binge-watching trends?”, the original responses encompass diverse perspectives, whereas the manipulated version exhibits a consistent bias aligned with the designated stance. These examples underscore Topic-FlipRAG’s capacity to impose coherent opinion shifts across multiple queries under a single topic.

### v3

@@ 4.3.2 Gradient-Based Optimization
Gradient Optimization: During each optimization step, we compute and combine gradients for all queries in $Q$. Inspired by [liu2022order], we introduce a word enhancement mechanism that incorporates query-specific keywords and lexical distributions derived from the model’s vocabulary. This guided selection of candidate tokens captures semantic overlaps among multiple queries and enhances overall trigger relevance.
@@
Beam Search: During the beam search phase, we evaluate candidate triggers, along with doc-specific augmentation, to select those that maximize combined similarity scores across the entire query set. This selection process ensures that the chosen trigger consistently enhances the ranking impact of the target passage when combined with $\text{doc}_{know}$.
@@
Doc-Specific Augmentation: At each iteration of gradient calculation and beam search, the intermediate soft representation of the trigger is evaluated only in the context of its concatenation with $\text{doc}_{know}$ and jointly across all queries. While this introduces additional computational overhead, it ensures tight contextual alignment with the target document, leading to more substantial gains in retrieval performance.
4.3.3 Final Output and RAG Opinion Influence
@@
After convergence, the adversarial trigger $T$ is appended to $\text{doc}_{know}$ to form the final adversarial document:
$\text{doc}_{adv}=[\text{doc}_{know},T].$
@@
This $\text{doc}_{adv}$ is then injected into the RAG system’s corpus $D^{\prime}$, effectively increasing its likelihood of being retrieved among the top-$k$ documents for each query $q\in Q$. Consequently, $\text{doc}_{adv}$ frequently appears in retrieval results, leading the LLM to generate outputs that align with the target stance $S_{t}$. This process effectively manipulates the overall stance of the RAG system’s output by ensuring that the augmented document is both prominently retrieved and influences the generated responses toward the target stance $S_{t}$.
5 Experiments
5.1 Research Questions
@@
We propose three research questions to evaluate the effectiveness of our method in the topic-queries task, focusing on black-box NRM attacks and opinion manipulation to RAGs.
@@
RQ1: Can Topic-FlipRAG effectively boost the rankings of target documents across queries in the retriever within RAG?
@@
5.2 Datasets
MSMARCO. We utilized the MS MARCO Passage Ranking dataset [nguyen2016ms] to evaluate the effectiveness of our method in improving document ranking in challenging topic-queries tasks. Specifically, we examined whether our method could effectively boost the retrieval rankings of target documents across queries in the retriever within RAG.
For topic-list construction, we applied K-means clustering to group similar queries, forming topics with related queries. To ensure robustness, we filtered topics based on intra-topic similarity, retaining only those with high semantic diversity and sufficient query count. This process resulted in 29 topics, each containing an average of 22.28 queries with a similarity score of approximately 0.5, thereby supporting a rigorous and diverse evaluation.
PROCON. We conducted our opinion manipulation experiments using a curated dataset of controversial topics sourced from the PROCON.ORG website. The dataset spans over 80 topics across diverse domains such as society, health, government, and education. Each topic is framed around two opposing stance labels {PRO (support), CON (oppose)}, with accompanying passages that present arguments from both perspectives.
To simulate real-world user interactions with a RAG system, we instructed GPT-4o to act as a proxy user and generate 40 candidate sub-queries for each topic. These sub-queries were crafted to reflect the diverse questions and concerns that users might pose when engaging with a specific controversial topic. To ensure semantic diversity, we applied a similarity-based filtering process, retaining only sub-queries with pairwise similarity scores below approximately 0.85. This step effectively removed redundancies while preserving a broad spectrum of perspectives. As a result, the final set of topic queries achieved an average similarity score of approximately 0.7, balancing semantic relevance with diversity. More detailed information on the datasets refer to the dataset description in our github repository [^2] https://github.com/LauJames/Topic-FlipRAG.
5.3 Experiment Details
(1) Black-box RAG. The black-box RAG process is denoted as $\text{RAG}_{\text{black}}$. The RAG framework is the Conversational RAG from LangChain for our experiments. The LLMs adopted in RAG are two widely adopted open-source instruction-tuned models: Meta-Llama-3.1-8B-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct (referred to as Llama3.1) and Qwen-2.5-7B-Instruct [^4] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct (Qwen2.5).
(2) Retrieval model. We benchmark three dominant dense retrievers, i.e., Contriever [gao2021unsupervised], DPR [karpukhin-etal-2020-dense], and ANCE[xiong2020approximate]. Following standard practice, we use dot product between the embedding vectors of a query/question and a candidate document as their similarity score $R$.
(3) Opinion classification. We adopt Qwen2.5-Instruct-72B as the opinion classifier. We selected high-performing LLMs that are widely recognized and adopted in the open-source community, based on the current GPU resources. The detailed description is provided in Appendix A.1
(4) Hyper-parameter settings. During the knowledge-guided attack process, we set the maximum editing distance $\epsilon$ to 0.2, the semantic similarity threshold $\lambda$ to 0.85, and the number of iterations $N$ to 5. For adversarial trigger generation, we use a beam size of 3, top-$k^{\prime}$ sampling with $k^{\prime}$ set to 10, a batch size of 32, a temperature of 1.0, a learning rate of 0.005, and a trigger sequence length of 10. In RAGblack configuration, the number of retrieved documents $K$ is set to 3, and the LLM temperature is fixed at 1.0.
(5) Target documents. For the PROCON dataset, we rank documents based on their relevance to each topic-query set $Q$ and target stance $S_{t}$, then select the five least relevant documents as poisoning targets. For MS MARCO, we first retrieve the top-1000 passages ranked by relevance per topic, then identify the passage with the lowest average rank across the corresponding queries. This selection strategy ensures that evaluation is conducted under challenging conditions by focusing on minimally relevant passages.
(6) Environment. All the methods run on a server configured with Python 3.8 environment, four NVIDIA DGX A100 GPUs (80 GB each), and 1 TB of system memory.
5.4 Baseline Settings
@@
To evaluate the effectiveness of our proposed method, we compare it against adversarial attack baselines tailored for black-box, topic-oriented RAG scenarios, with an emphasis on minimal modifications to the original documents. We exclude BadRAG[xue2024badrag], a backdoor RAG attack limited to white-box scenarios, as well as topic-IR-attack[liu2023topic], due to the lack of a complete implementation that prevents reliable reproduction. For the baseline methods, we adapt them to meet the requirements of our task while preserving their core components. A brief overview of these baselines is provided below:
@@
PoisonedRAG. Zou et al.[zou2024poisonedrag] propose an approach adaptable to both black-box and white-box settings. For our task, we employ its black-box strategy by inserting a randomly chosen query from the topic-queries set $Q$ at the beginning of each document.
@@
PAT. This gradient-based adversarial retrieval attack [liu2022order] uses a pairwise loss function to generate triggers that meet the fluency and coherence constraints. We adapt PAT to produce triggers $T_{\text{pat}}$ for target documents within the topic-queries set, evaluating their effectiveness under black-box conditions.
@@
Collision. This method generates adversarial paragraphs (collisions) via gradient-based optimization to produce content semantically aligned with the target query. In a topic-queries context, we create collisions for the entire topic-queries set and examine their transfer performance on the retriever within black-box RAG.
@@
These baseline methods provide benchmarks for comparing the efficacy of our approach in a black-box, topic-oriented RAG attack scenario.
5.5 Evaluation Metrics
For RQ1, which focuses on ranking manipulation, we report the change in top-3 target opinion proportion ($\text{Top3}_{\text{att}}-\text{Top3}_{\text{ori}}=\text{top3-v}$), Ranking Attack Success Rate (RASR), Boost Rank (BRank), and the percentage of target documents appearing in the Top-50 and Top-500. These metrics jointly reflect the effectiveness of rank promotion across queries.
top3-v. Computed by subtracting $\text{Top3}_{\text{ori}}$ from $\text{Top3}_{\text{att}}$, top3-v ranges from -1 to 1. A positive value signifies a successful increase of the target opinion in top-3 results, while a negative value indicates a detrimental effect.
Ranking Attack Success Rate (RASR). RASR captures how frequently target documents are successfully boosted in each query’s ranking.
Boost Rank (BRank). BRank is the average rank improvement for all target documents under each query.
@@ 5.5 Evaluation Metrics
Average Stance Variation (ASV). ASV is the absolute difference between manipulated and original opinion scores (i.e., 0 = opposing, 1 = neutral, 2 = supporting). Higher values indicate stronger polarity shifts and manipulation impact.
@@
6 Experimental Results Analysis
Table 1: White-box information retrieval (IR) attack performance (RASR, Top-50, and Top-500 are reported in %) on MS MARCO dataset. Bold indicates the best attack performance.
@@ 6 Experimental Results Analysis
Table 2: White-box and Black-box IR attack results (%) on PROCON dataset. Bold indicates the best attack performance.
@@
Table 3: Topic-Oriented RAG attack results on PROCON dataset, including ASV on different domains (Health & Environment, Government & Politics, Education, and Society & Culture), average ASV, and $\Delta$ ASV. Bold shows the best performances.
Method Target Stance Contriever + Qwen2.5 Contriever + Llama3.1
[…]
@@
6.1 RQ1: Can Topic-FlipRAG effectively boost the rankings of target documents across-queries in the retriever within RAG?
To explore RQ1, we first conducted white-box information retrieval (IR) manipulation experiments on MSMARCO and PROCON datasets. The results of these experiments are presented in Table 1 and Table 2, respectively.
MSMARCO Dataset. The results demonstrate that Topic-FlipRAG significantly outperforms all baselines across all evaluation metrics, even under challenging conditions where the MSMARCO dataset is configured with reduced intra-topic query similarity to simulate more extreme scenarios. As shown in Table 1, it achieves a near-perfect RASR of 99.29%, well above Poisoned-RAG (64.39%), Collision (93.19%), and PAT (89.63%). It also records the highest B-rank (521.30) and strong top-ranked document coverage (18.35% in top-50%, 79.49% in top-500%), demonstrating its effectiveness in simultaneously manipulating rankings across multiple queries.
PROCON Dataset. On the PROCON dataset (Table 2), the results also show the superiority of Topic-FlipRAG. Under white-box settings, for both target stances (CON and PRO), Topic-FlipRAG achieves the highest RASR (0.79 for CON, 0.75 for PRO) and the largest top3-v (0.45 for CON and 0.37 for PRO, respectively). In contrast, Poisoned-RAG, which uses a simple query+ strategy, shows considerably lower RASR (0.40 for CON and 0.33 for PRO, respectively), demonstrating the limitations of its approach for topic-oriented attacks. While the gradient-based methods, Collision and PAT, perform comparably to or even better than Poisoned-RAG in opinion manipulation, they still fall short of Topic-FlipRAG.
To further evaluate our method’s practical effectiveness, we conducted black-box IR manipulation experiments on the PROCON dataset using three widely adopted dense retrievers, e.g., Contriever, ANCE, DPR. Owing to the inherent robustness of these retrievers, performance declined for all attack methods compared to the white-box setting. Nevertheless, Topic-FlipRAG displays the most consistent performance, retaining relatively high manipulation effectiveness across all models and metrics. For instance, under Contriever, Topic-FlipRAG achieves the highest RASR (0.52 for CON and 0.49 for PRO, respectively) and top3-v (0.26 for CON and 0.22 for PRO, respectively), significantly outperforming Poisoned-RAG, Collision, and PAT. A similar performance pattern is observed with ANCE and DPR, where Topic-FlipRAG consistently maintains a clear advantage.
Topic-FlipRAG’s superior performance in transfer attack settings can be attributed to its integration of the LLM’s general semantic knowledge and NRM gradient guidance, along with a design specifically tailored for the topic-queries context. This combination enables the method to generate adversarial triggers that are both generalizable and contextually precise, allowing it to adapt effectively to varying retriever architectures without relying on the the parameters of the target IR model. In contrast, gradient-based methods like Collision and PAT depend heavily on gradient information from the target model, leading to substantial performance degradation in transfer attack scenarios. For instance, Collision achieves a maximum RASR of only 0.21 on Contriever for both CON and PRO stances, while PAT performs even worse, with RASR values of 0.16 (CON) and 0.13 (PRO).
@@
6.2 RQ2: To what extent does Topic-FlipRAG affect the answers generated by the target RAG systems?
Table 3 shows the opinion manipulation results on RAG outputs, comparing two RAG systems: Contriever + Qwen2.5 and Contriever + Llama3.1. More results with other retrievers (ANCE and DPR) are reported in Table 11 (Appendix).
[…]
The manipulation effectiveness of Topic-FlipRAG varies across topic domains. For instance, with Llama3.1, S&C questions are most prone to manipulation towards an opposing stance (avg-ASV = 0.55), whereas EDU questions are resistant to manipulation (avg-ASV = 0.41). It is more difficult to manipulate supporting stances in H&E questions than opposing ones. Both Qwen2.5 and Llama3.1 achieve higher avg-ASV scores for supporting stance manipulation in G&P, whereas EDU remains the most challenging domain.
In order to illustrate the manipulation results more vividly, we randomly selected the topic "Is Binge-Watching Good for You?" for a case study, as shown in Figure 10 (in Appendix). This topic comprises multiple sub-queries, enabling an examination of topic-level manipulation consistency. Before manipulation, responses span a mix of supportive, opposing, and neutral perspectives; after applying Topic-FlipRAG, the distribution shifts toward the target stance—becoming predominantly supportive under "Pro" (blue) or predominantly opposing under "Con" (red). For instance, in "Binge-watching: addictive or relaxing?", the original response presents a balanced view, while the manipulated versions adopt distinctly polarized stances—emphasizing relaxing effects under "Pro" and highlighting addictive harms under "Con." Similarly, for "Opinions on binge-watching trends?", the initially diverse responses become consistently biased toward the intended stance. These examples underscore Topic-FlipRAG’s ability to impose coherent opinion shifts across multiple queries within a single topic.